OpenStudy (anonymous):

If the probability of success during the 4th event of a geometric experiment is 0.0384, which of these could be the probability of success during the 5th event of the geometric experiment?

OpenStudy (perl):

Which of these?

OpenStudy (anonymous):

A.0.01536 B.0.24 C.0.6 D0.096

OpenStudy (perl):

Let p = probability of success Let q = 1 - p be the probability of failure. Then the probability of having the first success on the kth attempt, for k >= 1, is given by P(X = k) = q^(k - 1) * p

OpenStudy (perl):

probability of success during 4th event of a geometric experiment is 0.0384, which of these could be the probability of success during the 5th event of the geometric experiment? I get P( X=4) = q^(4-1) * p = .0384

OpenStudy (perl):

I think you can use the fact that since q and p are less than 1 , then the probability will get smaller on the next try

OpenStudy (perl):

q^3 * p = .0384 P(x=5) = q^4 * p , which must be smaller

OpenStudy (anonymous):

Yeah I thought so but thank you I understand much better.

OpenStudy (anonymous):

Thanks I got the answer it was A.